# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ 2. Lattice version of Halburd’s method
(2.1) $\ {\color[rgb]{1,0,0}x_{t-1,n+2}}\\ {\color[rgb]{1,0,0}x_{t-1,n+1}}\\ {\color[rgb]{1,0,0}x_{t-1,n}}&{\color[rgb]{1,0,0}x_{t,n}=\varepsilon}\\ &{\color[rgb]{1,0,0}x_{t,n-1}}&{\color[rgb]{1,0,0}x_{t+1,n-1}}&{\color[rgb]{1,0,0}x_{t+2,n-1}}&{\color[rgb]{1,0,0}\cdots}\end{matrix}$
@@
where “reg” means that the leading order with respect to $\varepsilon$ is 00 and the leading coefficient depends on some other initial variable.
@@
$\ {\color[rgb]{1,0,0}x_{0,4}}&x_{1,4}&x_{2,4}&x_{3,4}&x_{4,4}\\ {\color[rgb]{1,0,0}x_{0,3}}&x_{1,3}&x_{2,3}&x_{3,3}&x_{4,3}\\ {\color[rgb]{1,0,0}x_{0,2}}&x_{1,2}&x_{2,2}&x_{3,2}&x_{4,2}\\ {\color[rgb]{1,0,0}x_{0,1}}&{\color[rgb]{1,0,0}z}&x_{2,1}&x_{3,1}&x_{4,1}\\ &{\color[rgb]{1,0,0}x_{1,0}}&{\color[rgb]{1,0,0}x_{2,0}}&{\color[rgb]{1,0,0}x_{3,0}}&{\color[rgb]{1,0,0}x_{4,0}}&{\color[rgb]{1,0,0}\cdots}\end{matrix}.$
Each $x_{t,n}$ is expressed as a rational function in the initial values $x_{1,1}$, $x_{s,0}$ ($s\geq 1$) and $x_{0,m}$ ($m\geq 0$) and we focus on the degrees with respect to the initial variable $z=x_{1,1}$. As in Example 1.1, we suppose that the other initial values are generic and think of each $x_{t,n}$ as a rational function in $z$ as $x_{t,n}=x_{t,n}(z)\in\mathbb{C}(z)$. Our basic strategy is, by substituting various values $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ for $z$ as in the case of mappings, to verify all the possibilities for $x_{t,n}(z)$ to become 00 or $\infty$.
According to the above pattern, if $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ spontaneously satisfies $x_{t,n}(z^{*})=0$ for $t,n\geq 1$, then we have $x_{t+1,n}(z^{*})=\infty$, $x_{t,n+1}(z^{*})=\infty$ and $x_{t+1,n+1}(z^{*})=0$. Moreover, the equation has the following pattern II where $\infty$ appears as a value:
$\text{II}\colon\quad\begin{matrix}{\color[rgb]{1,0,0}\vdots}&&&&&\iddots\\ {\color[rgb]{1,0,0}\bullet}&\circ&\circ&\circ&\infty\\ {\color[rgb]{1,0,0}\bullet}&\circ&\circ&\infty&\circ\\ {\color[rgb]{1,0,0}\bullet}&\circ&\infty&\circ&\circ\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\infty}&\circ&\circ&\circ\\ &{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\cdots}\end{matrix}.$
@@
It is known that, in our settings, $\text{I}_{t,n}$ and II are the only possibilities where $\infty$ appears as a value. How to prove this, which we will discuss later, is one of the main topics in this paper.
@@
$\deg_{z}x_{t,n}(z)$ $=Z_{t,n}+Z_{t-1,n-1}$
$=Z_{t-1,n}+Z_{t,n-1}+\delta_{t,n}.$
The right-hand side of each line is the degree calculated as the number of preimages of 00 and $\infty$, respectively. In the above relations,
@@
The degrees calculated here are in perfect agreement with those calculated in [20].
@@
The second problem is how an equation enters a singularity. For example, when we calculated the singularity pattern starting with $x_{t,n}=0$ in the singularity confinement calculation, the initial values are taken as (2.1). However, we must confirm that the pattern starting with the spontaneous occurrence of $x_{t,n}(z^{*})=0$ in fact coincides with $\text{I}_{t,n}$, which is calculated by singularity confinement. The definition of the spontaneous occurrence of $x_{t,n}(z^{*})=0$ and its multiplicity should be clarified, too.
The third problem is the global structure of singularity patterns. Singularity confinement is a local property. For instance, when we calculated the singularity pattern starting with $x_{t,n}=0$ in the example, we did not discuss the outside region of (2.2). Therefore, singularity confinement does not guarantee that for any $z=z^{*}\in\mathbb{P}^{1}(\mathbb{C})$, an equation does not have any strange pattern. For example, we must confirm that two or more singularity blocks do not simultaneously appear (see Figures 1 and 2). This problem is slightly similar to how to distinguish open and periodic patterns, which we mentioned as a minor problem in the case of mappings. In the case of lattice equations, however, this is not a minor problem anymore because there are too many degrees of freedom in initial values. Moreover, we must show that each pattern has only one starting point (see Figure 3). Even in the case of Hirota’s discrete KdV equation, many types of complicated singularity patterns have been reported [44, 45]. Therefore, we must confirm that substituting a value only for $z$ does not generate such complicated patterns.
$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
Figure 1. Example of an undesirable singularity pattern, where a single value $z^{*}$ generates two or more independent singularity blocks. Such a phenomenon can never occur in the case of mappings.
$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
Figure 2. Similar to Figure 1, but after a singularity block is confined, another block appears. Such a phenomenon is common in the case of mappings since two or more patterns are sometimes combined and generate this type of pattern. A cyclic pattern of a mapping falls into this category, too.
$\times$$\times$$\times$$\times$$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
@@
The fourth problem is, as mentioned in the above example, how to confirm that there is no possibility for $x_{t,n}(z)$ to become 00 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II. For example, we must at least check that the spontaneous occurrence of $x_{t,n}(z^{*})=\beta\notin\{0,\infty\}$ does not generate 00 or $\infty$ as a value. This is the most difficult part of our theory and we must develop some techniques to resolve this problem.
The only remedy known today for these problems would be the coprimeness. This property was invented as an algebraic reinterpretation of singularity confinement and can deal with global singularity structure [26]. For example, in the case of Hirota’s discrete KdV equation, one can show that each iterate $x_{t,n}$ decomposes into
@@
where each $\tau_{t,n}$ is an irreducible polynomial with no monomial factor in the initial variables, $\rho_{t,n}$ is a monic Laurent monomial and $\tau_{t,n}$ and $\tau_{s,m}$ are coprime unless $(t,n)=(s,m)$. This decomposition, which obviously corresponds to the pattern $\text{I}_{t,n}$, tells us everything we need to know about singularities. For example, $x_{t,n}(z)$ should be defined to take a singularity at $z=z^{*}$ if $x_{t,n}(z^{*})\in\{0,\infty\}$. For $(t,n)\neq(1,1)$, the spontaneous occurrence of $x_{t,n}(z^{*})=0$ can be defined as $\tau_{t,n}(z^{*})=0$ ($z^{*}\neq 0,\infty$) and thus $Z_{t,n}$ is defined as $\deg_{z}\tau_{t,n}(z)$. Since each $\tau_{s,m}(z)$ is coprime to $\tau_{t,n}(z)$ unless $(s,m)=(t,n)$, such $z^{*}$ does not generate any 00 or $\infty$ other than the pattern $\text{I}_{t,n}$. The patterns $\text{I}_{1,1}$ and II correspond to $z=0$ and $\infty$, respectively. Therefore, the decomposition guarantees that there is no possibility for $x_{t,n}(z^{*})$ to become 00 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II.
@@
First, it requires too much computation to prove that an equation satisfies the coprimeness. Apparently, the coprimeness property is related to Laurentification, which is a technique to transform an equation into one with the Laurent property by transforming dependent variables [23, 26]. Although some researchers (the present author included) invented several techniques to show the coprimeness, no completely general method is known as of now. Therefore, in previous studies, some part of the discussion almost always relies on the concrete form of each equation [26, 27, 29, 25]. In addition, as seen in the above example, the theory of coprimeness is most compatible with singularity analysis when the singular values are 00 and $\infty$. If singularities take other values, we need multiple tau functions to express the original dependent variables, which makes Laurentification much more complicated [16, 17, 24].
In this paper, we do not use the theory of coprimeness to justify our calculations. Our basic strategy is based on the fact that an initial value problem for a lattice equation always has infinitely many initial variables. For example, consider the same domain as in Example 2.1 and suppose that $x_{t,n}(z^{*})=\alpha$ is a singularity. Then, on which initial variables does $z^{*}$ depend? Clearly, $z^{*}$ does not depend on $x_{t+1,0}$ or $x_{0,n+1}$ since $x_{t,n}(z)$ is independent of these initial variables. This idea, which definitely is not valid in the case of mappings, enables us to calculate all singularity patterns from basic computations such as (2.2).
@@ 3. How to calculate degrees for lattice equations
In this section, we state our main theorems (Theorems 3.9, 3.11, 3.12 and 3.15) and give some propositions, which is helpful to detect the singularity structure of an equation. All proofs will be given in the next section.
First, we define the conditions on a domain we consider in this paper.
@@
For a domain $H\subset\mathbb{Z}^{2}$, $H\neq\emptyset$, we consider the following two conditions:
@@
    $H_{\leq(t,n)}=\mathbb{Z}^{2}_{\leq(t,n)}\cap H=\{(t^{\prime},n^{\prime})\in H\mid t^{\prime}\leq t;\ n^{\prime}\leq n\},$
@@
    ($\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$
    is contained in $H$.
When considering a domain $H$ for a quad equation in this paper, we always assume that $H$ satisfies the above conditions. For a domain $H$, we define its initial boundary $H_{0}\subset H$ as
@@
One can check that $(t,n)\in H$ does not belong to the initial boundary if and only if $(t-1,n)$, $(t,n-1)$ and $(t-1,n-1)$ all belong to $H$.
The reason why we impose the above conditions when considering degree growth is discussed in [20].
From here on, we fix a quad equation (1.1), a domain $H\subset\mathbb{Z}$ and an arbitrary one of the initial variables $z=x_{t_{0},n_{0}}$. We think of each $x_{t,n}$ as a rational function in $z$, i.e., $x_{t,n}=x_{t,n}(z)$ and consider the individual degrees with respect to $z$. Note that among infinitely many choices of $H$ and $z$, only four of them are essential. We will leave such discussion in Appendix A (Theorem A.1) as it is only of secondary importance in this paper.
Definition 3.3.
@@
$\mathbb{K}=\mathbb{C}\left(x_{t^{\prime},n^{\prime}}\mid(t^{\prime},n^{\prime})\in H_{0};\ x_{t^{\prime},n^{\prime}}\neq z\right).$
@@
In the examples in §1 and §2, the initial variables other than $z$ are all taken generic. In our strategy, however, we do not use this approach. We think of each initial value other than $z$ just as a variable. Therefore, $x_{t,n}(z)$ is an element of $\mathbb{K}(z)$. Instead of taking generic initial values, we will substitute a $\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$-value $z^{*}$ for $z$. From here on, we simply use $\mathbb{P}^{1}$ to denote $\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$.
@@
For $(t,n)\in H$ and $z^{*}\in\mathbb{P}^{1}$, we say that $x_{t,n}(z)$ takes a constant singularity at $z=z^{*}$ if $x_{t,n}(z^{*})\in\mathbb{P}^{1}(\mathbb{C})$. The multiplicity of a constant singularity $x_{t,n}(z^{*})=x^{*}$ ($\in\mathbb{P}^{1}(\mathbb{C})$) is the multiplicity of $x_{t,n}(z)=x^{*}$ at $z=z^{*}$ as a rational function in $z$. We often think of the multiplicity of $x_{t,n}(z^{*})=x^{*}$ as 00 if $x_{t,n}(z^{*})\neq x^{*}$ (including the case where $x_{t,n}(z)$ does not take a constant singularity at $z=z^{*}$).
For $z^{*}\in\mathbb{P}^{1}$, the constant singularity pattern corresponding to $z=z^{*}$ consists of the following data:
  • the set of the points $(t,n)\in H$ where $x_{t,n}(z)$ takes a constant singularity at $z=z^{*}$,
  • their values, i.e., $x_{t,n}(z^{*})\in\mathbb{P}^{1}(\mathbb{C})$ for each point $(t,n)\in H$ where $x_{t,n}(z^{*})$ takes a constant singularity,
@@
We say that a constant singularity pattern corresponding to $z=z^{*}$ is fixed (resp. movable) if $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ (resp. $z^{*}\in\mathbb{P}^{1}\setminus\mathbb{P}^{1}(\mathbb{C})$). A constant singularity pattern is said to be confining (resp. non-confining) if it contains only finitely many (resp. infinitely many) constant singularities. We say that a constant singularity pattern is solitary if it consists of only one constant singularity.
Since all the singularity patterns we consider in the remaining part of this paper are constant singularity patterns, we sometimes use a singularity pattern (or simply a pattern) to denote a constant singularity pattern.
@@
We say that a quad equation $\Phi$ satisfies the $\partial$-factor condition if it satisfies the following two properties:
  • Let $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ be an arbitrary prime factor that appears in the numerator of $\partial_{B}\Phi$ or $\partial_{B}\frac{1}{\Phi}$. Then, $F$ satisfies one of the following:
@@
    That is, the numerators of $\partial_{B}\Phi$ and $\partial_{B}\frac{1}{\Phi}$ do not have a prime factor that depends on both $C$ and $D$ but not on $B$.
@@
We say that the equation satisfies the basic pattern condition for $x^{*}$ if any prime factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ of the above relation satisfies one of the following:
@@
That is, the equation does not satisfy the basic pattern condition for $x^{*}$ if the relation $\Phi(B,C,D)=x^{*}$ has a prime factor that contains both $B$ and $C$ but not $D$.

### v3

@@ 2. Lattice version of Halburd’s method
[…]
@@
where “reg” means that the leading order with respect to $\varepsilon$ is 0 and the leading coefficient depends on some other initial variable.
@@
$\begin{matrix}{\color[rgb]{1,0,0}\vdots}&&&&&\iddots\\ {\color[rgb]{1,0,0}x_{0,4}}&x_{1,4}&x_{2,4}&x_{3,4}&x_{4,4}\\ {\color[rgb]{1,0,0}x_{0,3}}&x_{1,3}&x_{2,3}&x_{3,3}&x_{4,3}\\ {\color[rgb]{1,0,0}x_{0,2}}&x_{1,2}&x_{2,2}&x_{3,2}&x_{4,2}\\ {\color[rgb]{1,0,0}x_{0,1}}&{\color[rgb]{1,0,0}z}&x_{2,1}&x_{3,1}&x_{4,1}\\ &{\color[rgb]{1,0,0}x_{1,0}}&{\color[rgb]{1,0,0}x_{2,0}}&{\color[rgb]{1,0,0}x_{3,0}}&{\color[rgb]{1,0,0}x_{4,0}}&{\color[rgb]{1,0,0}\cdots}\end{matrix}.$
Each $x_{t,n}$ is expressed as a rational function in the initial values $x_{1,1}$, $x_{s,0}$ ($s\geq 1$) and $x_{0,m}$ ($m\geq 0$) and we focus on the degrees with respect to the initial variable $z=x_{1,1}$. As in Example 1.1, we suppose that the other initial values are generic and think of each $x_{t,n}$ as a rational function in $z$ as $x_{t,n}=x_{t,n}(z)\in\mathbb{C}(z)$. Our basic strategy is, by substituting various values $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ for $z$ as in the case of mappings, to verify all the possibilities for $x_{t,n}(z)$ to become 0 or $\infty$.
As in Example 1.1, we consider the “spontaneous occurrence” of $x_{t,n}(z^{*})=\alpha$. Although its mathematical definition will be given in Definition 4.14, we provide here a temporary interpretation: we say that $x_{t,n}(z^{*})$ spontaneously becomes $\alpha$ if $x_{t,n}(z^{*})=\alpha$ and $x_{s,m}(z^{*})$ is generic for all $(s,m)\leq(t,n)$ with $(s,m)\neq(t,n)$.
According to the above pattern, if $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ makes $x_{t,n}(z^{*})$ spontaneously become 0 for $t,n\geq 1$, then we have $x_{t+1,n}(z^{*})=\infty$, $x_{t,n+1}(z^{*})=\infty$ and $x_{t+1,n+1}(z^{*})=0$. In addition to the pattern $I_{t,n}$, the equation has the following pattern II where $\infty$ appears as a value:
[…]
@@
It is known that, in our settings, $\text{I}_{t,n}$ and II are the only possibilities where $\infty$ and 0 appear as a value, whereas in other settings more possibilities may occur (see [49, 50]). How to prove this, which we will discuss later, is one of the main topics in this paper.
@@
$\deg_{z}x_{t,n}(z)$ $=Z_{t,n}+Z_{t-1,n-1}$ $(\text{$\#$ of preimages of $0$})$
$=Z_{t-1,n}+Z_{t,n-1}+\delta_{t,n}$ $(\text{$\#$ of preimages of $\infty$}),$
where $\delta_{t,n}$ denotes Kronecker’s delta. In the above relations,
@@
The degrees calculated here are in perfect agreement with those calculated in [22], and the exact solutions for $Z_{t,n}$ and $\deg_{z}x_{t,n}(z)$ are
$Z_{t,n}=\min(t,n),\quad\deg_{z}x_{t,n}(z)=2\min(t,n)-1.$
@@
The second problem is how an equation enters a singularity. For example, when we calculated the singularity pattern starting with $x_{t,n}=0$ in the singularity confinement calculation, the initial values are taken as (2.1). However, we must confirm that the pattern starting with a spontaneous occurrence of $x_{t,n}(z^{*})=0$ in fact coincides with $\text{I}_{t,n}$, which is calculated by singularity confinement. The definition of the spontaneous occurrence of $x_{t,n}(z^{*})=0$ and its multiplicity should be clarified, too.
The third problem is the global structure of singularity patterns. Singularity confinement is a local property. For instance, when we calculated the singularity pattern starting with $x_{t,n}=0$ in the example, we did not discuss the outside region of (2.2). Therefore, singularity confinement does not guarantee that for any $z=z^{*}\in\mathbb{P}^{1}(\mathbb{C})$, an equation does not have any strange pattern. For example, we must confirm that one singularity pattern does not have two or more singularity blocks (see Figures 1 and 2). Here, a singularity block denotes a maximal subset of the singular part in which any two points can be joined by a finite sequence of steps of Euclidean length $1$ or $\sqrt{2}$. This problem is slightly similar to how to distinguish open and periodic patterns, which we mentioned as a minor problem in the case of mappings. In the case of lattice equations, however, this is not a minor problem anymore because there are too many degrees of freedom in initial values. Moreover, we must show that each pattern has only one starting point (see Figure 3). Even in the case of Hirota’s discrete KdV equation, many types of complicated singularity patterns have been reported [49, 50]. Therefore, we must verify that assigning a value to $z$, while keeping all other initial values generic, does not lead to such complicated singularity patterns.
[…]
Figure 1. Example of an undesirable singularity pattern, where a single value $z^{*}$ generates two or more independent singularity blocks. Here, we call two singularity blocks independent if they cannot be compared under the product order on $\mathbb{Z}^{2}$. Such a phenomenon can never occur in the case of mappings, since all points are totally ordered on $\mathbb{Z}$.
[…]
Figure 2. Similar to Figure 1, but two singularity blocks are not independent, as after a singularity block is confined, another block appears. Such a phenomenon is common in the case of mappings since two or more patterns are sometimes combined and generate this type of pattern. A cyclic pattern of a mapping falls into this category, too.
[…]
@@
The fourth problem is, as mentioned in the above example, how to confirm that there is no possibility for $x_{t,n}(z)$ to become 0 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II. For example, we must at least check that a spontaneous occurrence of $x_{t,n}(z^{*})=\beta\notin\{0,\infty\}$ does not generate 0 or $\infty$ as a value. As illustrated in Figure 2, a pattern starting with $\beta_{1}\notin\{0,\infty\}$ and later involving $\beta_{2}=0$ would be problematic. We must ensure that such a situation does not occur. This is the most difficult part of our theory and we must develop some techniques to resolve this problem.
The only remedy known today for these problems would be the coprimeness. This approach was developed as an algebraic reinterpretation of singularity confinement and can deal with global singularity structure [29]. For example, in the case of Hirota’s discrete KdV equation, one can show that each iterate $x_{t,n}$ decomposes into
@@
where $\tau_{t,n}$ is an irreducible polynomial with no monomial factor in the initial variables, $\rho_{t,n}$ is a monic Laurent monomial, and $\tau_{t,n}$ and $\tau_{s,m}$ are coprime unless $(t,n)=(s,m)$. This decomposition, which corresponds to the pattern $\text{I}_{t,n}$, tells us everything we need to know about singularities. In this illustrative example, $x_{t,n}(z)$ is considered to have a singularity at $z=z^{*}$ if $x_{t,n}(z^{*})\in\{0,\infty\}$; this convention applies only here, and the general definition will appear in Definition 3.5. For $(t,n)\neq(1,1)$, the spontaneous occurrence of $x_{t,n}(z^{*})=0$ can be defined as $\tau_{t,n}(z^{*})=0$ ($z^{*}\neq 0,\infty$) and thus $Z_{t,n}$ is defined as $\deg_{z}\tau_{t,n}(z)$. Since each $\tau_{s,m}(z)$ is coprime to $\tau_{t,n}(z)$ unless $(s,m)=(t,n)$, such $z^{*}$ does not generate any 0 or $\infty$ other than the pattern $\text{I}_{t,n}$. The patterns $\text{I}_{1,1}$ and II correspond to $z=0$ and $\infty$, respectively. Therefore, the decomposition guarantees that there is no possibility for $x_{t,n}(z^{*})$ to become 0 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II.
@@
First, it requires too much computation to prove that an equation satisfies the coprimeness. Apparently, the coprimeness property is related to Laurentification, which is a technique to transform an equation into one with the Laurent property by transforming dependent variables [25, 29]. Although some researchers (the present author included) invented several techniques to show the coprimeness, no general method is known as of now. Therefore, in previous studies, some part of the discussion almost always relies on the concrete form of each equation [29, 30, 32, 28]. In addition, as seen in the above example, the theory of coprimeness is most compatible with singularity analysis when the singular values are 0 and $\infty$. If singularities take other values, we need multiple tau functions to express the original dependent variables, which makes Laurentification much more complicated [18, 19, 26].
In this paper, we do not use the theory of coprimeness to justify our calculations. Our approach is based on the fact that an initial value problem for a lattice equation has infinitely many initial variables. For example, consider the same domain as in Example 2.1 and suppose that $x_{t,n}(z^{*})=\alpha$ is a singularity. Then, on which initial variables does $z^{*}$ depend? Clearly, $z^{*}$ does not depend on $x_{t+1,0}$ or $x_{0,n+1}$ since $x_{t,n}(z)$ is independent of these initial variables. This independence, which is not valid in the case of mappings, enables us to calculate all singularity patterns from basic computations such as (2.2).
@@ 3. How to calculate degrees for lattice equations
In this section, we state our main theorems (Theorems 3.14, 3.16, 3.17 and 3.20) and give some propositions, which is helpful to detect the singularity structure of an equation. All proofs will be given in the next section.
First, we define the past and future light cones.
Definition 3.1 (Past and Future Light Cones).
  1. (1)
    For $(t,n)\in\mathbb{Z}^{2}$, we define the past light cone emanating from $(t,n)$ as
    $\mathbb{Z}^{2}_{\leq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\leq t;\ n^{\prime}\leq n\}.$
  2. (2)
    For $(t,n)\in\mathbb{Z}^{2}$, we define the future light cone emanating from $(t,n)$ as
    $\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}.$
Next, we define the conditions on a domain we consider in this paper.
@@
  1. (1)
    A nonempty set $H\subset\mathbb{Z}^{2}$ is called a domain if the following two conditions are satisfied:
@@
      $H_{\leq(t,n)}:=\mathbb{Z}^{2}_{\leq(t,n)}\cap H=\{(t^{\prime},n^{\prime})\in H\mid t^{\prime}\leq t;\ n^{\prime}\leq n\},$
@@
      $\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$
      is contained in $H$. In particular, $H_{\geq(t,n)}:=\mathbb{Z}^{2}_{\geq(t,n)}\cap H$ coincides with $\mathbb{Z}^{2}_{\geq(t,n)}$.
  2. (2)
    For a domain $H$, we define its initial boundary $H_{0}\subset H$ as
@@
One can check that $(t,n)\in H$ does not belong to the initial boundary if and only if (t1,n)(t-1,n), (t,n1)(t,n-1) and (t1,n1)(t-1,n-1) all belong to $H$.
The reason why we impose the above conditions when considering degree growth is discussed in [22]. Domains satisfying these conditions were termed light-cone regular domains in [37].
From here on, we fix a quad equation (1.2), a domain $H\subset\mathbb{Z}$ and one of the initial variables $z=x_{t_{0},n_{0}}$. We think of each $x_{t,n}$ as a rational function in $z$, i.e., $x_{t,n}=x_{t,n}(z)$ and consider the degree with respect to $z$. Note that among infinitely many choices of $H$ (a domain) and $z$ (one of the initial variables), only four of them are essential, in the sense that for any choice of $H$ and $z$, the degree computation can be reduced to one of these four cases. We will leave such discussion in Appendix A (Theorem A.1) as it is only of secondary importance in this paper.
Definition 3.4 ($\mathbb{K}$).
@@
$\mathbb{K}=\mathbb{C}\left(x_{t,n}\mid(t,n)\in H_{0};\ x_{t,n}\neq z\right).$
@@
In the examples in §1 and §2, the initial variables other than $z$ are all taken generic. In our strategy, however, we do not use this approach. We think of each initial value other than $z$ just as a variable. Therefore, $x_{t,n}(z)$ is an element of $\mathbb{K}(z)$. Instead of taking generic initial values, we will substitute a $\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$-value $z^{*}$ for $z$, where $\overline{\mathbb{K}}$ is the algebraic closure of $\mathbb{K}$.
@@
For $(t,n)\in H$ and $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$, we say that $x_{t,n}(z)$ has a constant singularity at $z=z^{*}$ if $x_{t,n}(z^{*})\in\mathbb{P}^{1}(\mathbb{C})$. The multiplicity of a constant singularity $x_{t,n}(z^{*})=x^{*}$ ($\in\mathbb{P}^{1}(\mathbb{C})$) is the multiplicity of $x_{t,n}(z)=x^{*}$ at $z=z^{*}$ as a rational function in $z$. We define the multiplicity of $x_{t,n}(z^{*})=x^{*}$ as 0 if $x_{t,n}(z^{*})\neq x^{*}$ (including the case where $x_{t,n}(z)$ does not have a constant singularity at $z=z^{*}$). Here, a constant singularity $x_{t,n}(z^{*})=x^{*}$ consists of its position $(t,n)$, the corresponding value $x^{*}$, and its multiplicity. As in Example 1.1, when we are interested only in the value of a constant singularity, we use the term “singular value.”
For $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$, the constant singularity pattern corresponding to $z=z^{*}$ consists of the following data:
  • the set of the points $(t,n)\in H$ where $x_{t,n}(z)$ has a constant singularity at $z=z^{*}$,
  • their values, i.e., $x_{t,n}(z^{*})\in\mathbb{P}^{1}(\mathbb{C})$ for each point $(t,n)\in H$ where $x_{t,n}(z^{*})$ has a constant singularity,
@@
We say that a constant singularity pattern corresponding to $z=z^{*}$ is fixed (resp. movable) if $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ (resp. $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)\setminus\mathbb{P}^{1}(\mathbb{C})$). A constant singularity pattern is said to be confining (resp. non-confining) if it contains only finitely many (resp. infinitely many) constant singularities. We say that a constant singularity pattern is solitary if it consists of only one constant singularity.
Since all the singularity patterns we consider in the remaining part of this paper are constant singularity patterns, we sometimes use a singularity pattern (or simply a pattern) to denote a constant singularity pattern. We will discuss non-constant singularity patterns in §6.
Remark 3.6.
In the context of singularity confinement, a singularity is defined by loss of information on the initial values, as seen in Example 1.1. However, the definition of a constant singularity does not involve any loss of information on the initial values, although such a loss may occur in significant examples. If we were to require such a loss as part of the definition, it would be necessary to define what loss of information means for lattice equations. In fact, at the stage where the first singularity appears, no loss of information on the initial values seems to occur. This observation will later be confirmed in the proof of Theorem 3.16, although the theorem itself serves a different purpose. As this paper does not aim to establish an ideal definition of singularity confinement for lattice equations, we will not pursue this discussion further.
@@
Remark 3.8.
Note that for generic $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$, the corresponding basic pattern is solitary. That is, there are only finitely many values of $x^{*}$ that produce a non-solitary pattern. This fact will be justified later (see Proposition 3.24).
@@
We say that a quad equation $\Phi$ satisfies the $\partial$-factor condition if it has the following two properties:
  • Consider the factorizations of the numerators of $\partial_{B}\Phi$ and $\partial_{B}\frac{1}{\Phi}$, where $\partial_{X}=\frac{\partial}{\partial X}$ denotes the (partial) derivative with respect to a variable $X$. Let $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ be an arbitrary irreducible polynomial occurring in one of these factorizations. Then, $F$ satisfies one of the following:
@@
    That is, the numerators of $\partial_{B}\Phi$ and $\partial_{B}\frac{1}{\Phi}$ do not have an irreducible factor that depends on both $C$ and $D$ but not on $B$.
@@
Definition 3.10.
Let $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ and let
$\Phi(B,C,D)=\frac{\Phi_{1}(B,C,D)}{\Phi_{2}(B,C,D)}$
where $\Phi_{1},\Phi_{2}\in\mathbb{C}[B,C,D]$ are coprime. If $x^{*}\neq\infty$ (resp. $x^{*}=\infty$), an irreducible factor of the relation $\Phi(B,C,D)=x^{*}$ is an irreducible factor of the polynomial
$\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$
(resp. of the polynomial $\Phi_{2}(B,C,D)$).
@@
We say that the equation satisfies the basic pattern condition for $x^{*}$ if any irreducible factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ of the above relation satisfies one of the following:
@@
That is, the equation does not satisfy the basic pattern condition for $x^{*}$ if the relation $\Phi(B,C,D)=x^{*}$ has an irreducible factor that contains both $B$ and $C$ but not $D$.